1. Find the probability of 7 tomatos getting accepting in 6 stores when the mean is 5.5
2.Random sample of 120 customers’s spent an average of 9.2 hours on their professional job with a sample standard deviation of 5.1 hours. Calculate the specification hours with confidence of 95%.
3.In a set of 10 samples of bike, the mileage were listed as follows
41, 42, 45, 42, 47, 59, 55, 60, 40, 35
What percent of samples are getting more than 40 mileages?

Answer by stanbon(48569)

1. Find the probability of 7 tomatos getting accepting in 6 stores when the mean is 5.5
---
Mean of what? Do you have a standard deviation?
--------------------

2.Random sample of 120 customers’s spent an average of 9.2 hours on their professional job with a sample standard deviation of 5.1 hours. Calculate the specification hours with confidence of 95%.
---------------
x-bar = 9.2 hrs
E = 1.96*(5.1/sqrt(120)) = 0.9125
----
95% CI: 9.2 - 0.9125 < u < 9.2 + 0.9125
========================================

3.In a set of 10 samples of bike, the mileage were listed as follows
41, 42, 45, 42, 47, 59, 55, 60, 40, 35
What percent of samples are getting more than an average of 40 ?
---------
# above 40 = 8
Total number = 10
-----------------------
Percentage = 8/10 = 80%
=============================
Cheers,
Stan H.